# ToolACE-MT: Non-Autoregressive Generation for Agentic Multi-Turn Interaction

### v1

@@ Introduction
In this subsection, we compare the cost and quality of generating agentic dialogue data using ToolACE-MT versus MAS. As shown in Table 4, MAS yields a lower Offline Verification pass rate ( vs. ), therefore requiring a larger initial dataset and in total 275k API calls, to obtain 8,000 valid samples, significantly more than ToolACE-MT (188k). Models trained on ToolACE-MT data also perform better, demonstrating both higher efficiency and effectiveness.
@@
We further test ToolACE-MT with GPT-4o-mini, which results in a much lower pass rate and increased API calls, due to more frequent formatting errors and hallucinations. This reinforces that generating long, tool-intensive dialogues demands strong long-context handling, which smaller models like GPT-4o-mini and LLaMA3.1-8B-Inst (in our attempt, it failed to produce valid instances in most of time thus cannot generate sufficient usable data for training) struggle with.
@@
Finally, even after filtering, the model trained on GPT-4o-mini generated data still show a notable performance gap compared to that trained on GPT-4o generated data ( vs. ), highlighting that initial generation quality remains crucial despite post-processing.
Task Completion Efficiency.
@@
In addition to generation efficiency, we further examine how our non-autoregressive pipeline influences task completion efficiency. We hypothesize that this generation paradigm supports more effective overall task planning, thereby reducing the number of interaction turns required to complete a task. In contrast, MAS often involves trial-and-error behavior from the assistant model to identify correct actions. This hypothesis is first supported by statistics from the training data: as shown in the left part of Figure 4, instances generated by our method ToolACE-MT have fewer assistant turns on average than those generated by MAS. Evaluation on -Bench (right part) further validates this advantage, where our model completes tasks successfully with an average of assistant turns, compared to turns for MAS. These findings suggest that ToolACE-MT leads to better task structuring and more efficient interaction patterns.
Data Effectiveness and Generalizability
Iterative Refinement Time Scaling.
@@
In the previous subsection, we mentioned the complementary roles of Iterative Refinement and Offline Verification, both of which contribute to enhancing final data quality. To further investigate their interaction, we conduct an experiment where we vary the number of Iterative Refinement steps, specifically by applying more Reasonability Refinement operations (as Complexity Injection is not well-suited for repeated application within a single dialogue). For each refinement level, we train two models: one using data that has passed Offline Verification and one without. The performance trends are illustrated in Figure 5.
@@
Different Backbones.
To evaluate the generalizability of our generated data across different backbone models, we conduct experiments using base models of similar sizes, including Qwen2.5-7B-Inst and Qwen3-8B (no-thinking). The results, shown in Figure 6, include comparisons between raw models (without training), models trained with MAS data, and models trained with ToolACE-MT. As observed, both backbones benefit from training with MAS data, and training with ToolACE-MT leads to further consistent gains.
@@
Model Size Scaling.
Scaling laws suggest a strong correlation between model size and performance. To explore the scalability of function calling capabilities after training on our generated data, we evaluate the Qwen-2.5-xB-Inst series across a range of model sizes (0.5B, 1.5B, 3B, and 7B). Both the raw and fine-tuned versions (trained on our generated 8000 instances) are assessed on BFCL-v3, with results (Multi-Turn and overall) shown in Figure 7. As expected, larger models consistently outperform smaller ones. The smaller raw models (0.5B and 1.5B) exhibit little to no multi-turn capabilities, but fine-tuning with our dataset can enhance the corresponding performance. Notably, the improvements are more pronounced in the 3B and 7B models, suggesting that multi-turn function calling remains a relatively advanced ability that small models struggle to acquire. Overall, the fine-tuned models demonstrate a clear scaling trend, reinforcing the effectiveness of our data in equipping larger LLMs with complex function calling skills.
Conclusion

### v2

@@ 4.4 Data Efficiency
Data Generation Efficiency. We compare the cost and quality of generating agentic dialogue data using ToolACE-MT versus MAS. As shown in Table 5, MAS yields a lower Offline Verification pass rate ( vs. ), therefore requiring a larger initial dataset and in total 275k API calls, to obtain 8,000 valid samples, significantly more than ToolACE-MT (188k). To clarify token usage, we report total input/output tokens and the corresponding cost (GPT-4o pricing: /M input, /M output). The main efficiency gain comes from reduced input tokens, since fewer iterations are needed to generate longer multi-turn dialogues. Models trained on ToolACE-MT data also perform better, demonstrating both higher efficiency and effectiveness.
@@
Model Choices. We further test ToolACE-MT with GPT-4o-mini (results also shown in Table 5), which results in a much lower pass rate and increased API calls, due to more frequent formatting errors and hallucinations. This reinforces that generating long, tool-intensive dialogues demands strong long-context handling, which smaller models like GPT-4o-mini and LLaMA3.1-8B-Inst (in our attempt, it failed to produce valid instances in most of time thus cannot generate sufficient usable data for training) struggle with.
@@
Finally, even after filtering, the model trained on GPT-4o-mini generated data still show a notable performance gap compared to that trained on GPT-4o generated data ( vs. ), highlighting that generation quality before verification remains crucial despite post-processing.
@@
Task Completion Efficiency. In addition to generation efficiency, we further examine how our non-autoregressive pipeline influences task completion efficiency. Here we define task completion efficiency as how effectively the assistant completes a user task, where turn numbers per task serve as a signal. We hypothesize that this generation paradigm supports more effective overall task planning, thereby reducing the number of interaction turns required to complete a task. In contrast, MAS often involves trial-and-error behavior from the assistant model to identify correct actions. This hypothesis is first supported by statistics from the training data: as shown in the left part of Figure 5, instances generated by our method ToolACE-MT have fewer assistant turns on average than those generated by MAS. Evaluation on -Bench (right part) further validates this advantage, where our model completes tasks successfully with an average of assistant turns, compared to turns for MAS. These findings suggest that ToolACE-MT leads to better task structuring and more efficient interaction patterns.
4.5 Data Effectiveness and Generalizability
@@
Iterative Refinement Time Scaling. In the previous subsection, we mentioned the complementary roles of Iterative Refinement and Offline Verification, both of which contribute to enhancing final data quality. To further investigate their interaction, we conduct an experiment where we vary the number of Iterative Refinement steps, specifically by applying more Reasonability Refinement operations (as Complexity Injection is not well-suited for repeated application within a single dialogue). For each refinement level, we train two models: one using data that has passed Offline Verification and one without. The performance trends are illustrated in Figure 5.
@@
Different Backbones. To evaluate the generalizability of our generated data across different backbone models, we conduct experiments using base models of similar sizes, including Qwen2.5-7B-Inst and Qwen3-8B (no-thinking). The results, shown in Figure 7, include comparisons between raw models (without training), models trained with MAS data, and models trained with ToolACE-MT. As observed, both backbones benefit from training with MAS data, and training with ToolACE-MT leads to further consistent gains.
@@
Model Size Scaling. Scaling laws suggest a strong correlation between model size and performance. To explore the scalability of function calling capabilities after training on our generated data, we evaluate the Qwen-2.5-xB-Inst series across a range of model sizes (0.5B, 1.5B, 3B, and 7B). Both the raw and fine-tuned versions (trained on our generated 8000 instances) are assessed on BFCL-v3, with results (Multi-Turn and overall) shown in Figure 7. As expected, larger models consistently outperform smaller ones. The smaller raw models (0.5B and 1.5B) exhibit little to no multi-turn capabilities, but fine-tuning with our dataset can enhance the corresponding performance. Notably, the improvements are more pronounced in the 3B and 7B models, suggesting that multi-turn function calling remains a relatively advanced ability that small models struggle to acquire. Overall, the fine-tuned models demonstrate a clear scaling trend, reinforcing the effectiveness of our data in equipping larger LLMs with complex function calling skills.
5 Conclusion
@@
Appendix A More Training Details
Table 6 displays the selections of hyper-parameter, where all are selected based on performance of development set. All the models are trained with V100 8 based on LLaMA-Factory222https://github.com/hiyouga/LLaMA-Factory.
| Hyper-parameters | Selection Ranges | Final Selction |
| Learning rate | [5e-5, 1e-4, 2e-4] | 1e-4 |
| Batch size | [32, 64] | 64 |
| Warmup ratio | [0.1] | 0.1 |
| Lora rank | [8, 16] | 16 |
| Lora alpha | [16, 32] | 32 |
Appendix B Use of LLMs
We used large language models (LLMs) solely for writing assistance, specifically to polish the language of the manuscript. The models were employed to improve readability, grammar, and style, without altering the technical content, analysis, or results. All ideas, experimental designs, and conclusions were conceived and developed by the authors.
Appendix C Full Results on ACEBench
Full results on ACEBench are shown in Table 7, where additional results for single-turn categories are added (“S-Turn” is short for Single-Turn, “M-Turn” is short for Multi-Turn). The results reflect that ToolACE-MT still outperforms baselines in those single-turn categories, with a higer overall result.
| Models | Atom | S-Turn | M-Turn | Similar | Preference | Special | Agent (EA) | Agent (PA) | Overall |
| GPT-4o-2024-11-20 | 90.0 | 78.0 | 68.0 | 80.0 | 78.0 | 92.7 | 56.0 | 77.8 | 81.1 |
| Llama3.1-70B-Inst | 83.7 | 71.5 | 61.0 | 74.0 | 66.0 | 29.3 | 41.0 | 62.5 | 57.9 |
| Llama3.1-8B-Inst | 52.7 | 45.0 | 24.0 | 54.0 | 50.0 | 9.3 | 6.7 | 18.3 | 30.4 |
| Multi-Agent Simulation | 81.3 | 63.5 | 48.0 | 70.0 | 64.0 | 5.3 | 6.7 | 15.0 | 43.8 |
| ToolACE-MT | 83.0 | 64.0 | 51.0 | 68.0 | 68.0 | 8.7 | 8.4 | 34.0 | 45.2 |
| - Offline Verification | 77.1 | 61.0 | 44.0 | 60.0 | 64.0 | 8.7 | 1.7 | 28.5 | 41.8 |
| - Iterative Refinement | 61.7 | 56.0 | 34.0 | 56.0 | 50.0 | 5.3 | 1.7 | 22.8 | 34.5 |
Appendix D Additional Experiments
To further highlight the value of our generated data, we conducted additional experiments on three representative open-source datasets (ToolACE-10k333https://huggingface.co/datasets/Team-ACE/ToolACE, Toucan-1.5M444https://huggingface.co/datasets/Agent-Ark/Toucan-1.5M, ToolDial-9k555https://github.com/holi-lab/ToolDial/tree/main), where we train several models using the same training setting and the corresponding results on the three benchmarks are summarized in Table 8. As shown, except for the -Airline dataset (for which we previously noted reliability issues), our ToolACE-MT consistently achieves superior performance across all multi-turn evaluations. Although it lags slightly behind ToolACE (which primarily contains single-turn samples) in the single-turn setting, ToolACE-MT still delivers substantial improvements over the base Llama3.1-8B model. Overall, the results indicate that the data generated by ToolACE-MT matches or surpasses that produced by existing methods, with particularly strong gains on multi-turn scenarios, underscoring the effectiveness and superiority of our ToolACE-MT approach.
| Models | BFCL | ACEBench | -Bench | |||
| Single-Turn | Multi-Turn | Multi-Turn | Agent-PA | -Retail | -Airline | |
| Llama3.1-8B-Inst | 72.7 | 9.3 | 24.0 | 18.3 | 6.1 | 26.0* |
| + ToolACE-10k | 80.0 | 8.6 | 51.0 | 14.0 | 13.9 | 22.0 |
| + Toucan-1.5M (sampled 8k) | 75.2 | 15.0 | 41.0 | 22.6 | 18.3 | 18.0 |
| + ToolDial-9k | 70.1 | 12.3 | 30.0 | 11.4 | 10.4 | 20.0 |
| + MAS (this work) | 79.2 | 31.4 | 48.0 | 15.0 | 21.7 | 10.0 |
| + ToolACE-MT (this work) | 78.2 | 40.3 | 51.0 | 34.0 | 25.2 | 16.0 |
Appendix E Failure Cases Analysis
This section analyzes the most frequent failure modes encountered during Offline Verification. Table 9 summarizes the top three failure types for both rule-based and model-based verification. As shown, most errors stem from parameter-related issues, including incorrect parameter values, mismatches between parameters and function signatures, and improper or incomplete parameter formatting. These findings highlight that parameter handling remains the primary bottleneck in achieving fully reliable tool-calling data generation.
| Rule-based Problems | Ratio | Model-based Problems | Ratio |
| Parameter Hallucination | 51% | Parameter Extraction | 23% |
| No Tool Calling | 18% | Tool Chosen/Ordering | 21% |
| Conversation Role Misordering | 11% | Parameter Format Compliance | 16% |
Appendix F Data Example
Below we show a data example for reference.
[system] You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/ tool calls to achieve the purpose. If none of the functions can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the function calls in your response. If you decide to invoke any of the function(s), you MUST put it in the format of [func_name1(params_name1=params_value1, params_name2=params_value2…), func_name2(params), …] You SHOULD NOT include any other text in the response. Here is a list of functions in JSON format that you can invoke:
{”name”: ”get_curr_date”, ”description”: ”Retrieves the current date in YYYY-MM-DD format.”, ”parameters”: …}
{”name”: ”get_curr_weekday”, ”description”: ”Retrieves the current day of the week.”, ”parameters”: …}
{”name”: ”book_flight”, ”description”: ”Books a flight between two cities on a given date.”, ”parameters”: …}
[user] I’m planning a trip and need to fly to Shanghai. Could you help me book a flight for next Tuesday?
[assistant] Could you provide your departure city?
[user] Singapore.
[assistant] [get_curr_date(), get_curr_weekday()]
[tool] [{”current_date”: ”2023-10-05”}, {”current_ weekday”: ”Thursday”}]
[assistant] [book_flight(departure_city=’Singapore’, arrival_city=’Shanghai’, departure_date=’2023-10-10’)]
[tool] [{”confirmation”: ”Flight successfully booked from Singapore to Shanghai on 2023-10-10.”}]
[assistant] Great news! Your flight from Singapore to Shanghai for next Tuesday, October 10th, has been successfully booked. Feel free to let me know if there’s anything else you need assistance with!
[...] … (More Turns)
The data highlights several essential abilities that an agent must possess. For example:
1) Asking clarification questions when user input is ambiguous or incomplete.
2) Multi-step function calling by leveraging implicit cues from historical context. Here the assistant needs to book a flight, but the book_flight tool requires a specific date in YYYY-MM-DD format. As the date is not provided explicitly, the agent should infer it using supporting tools including get_curr_date and get_curr_weekday.